# The Power of Graph Sparsification in the Continual Release Model

### v1

@@ Abstract
The graph continual release model of differential privacy seeks to produce differentially private solutions to graph problems under a stream of updates where new private solutions are released after each update. Streaming graph algorithms in the non-private literature also produce (approximately) accurate solutions when provided updates in a stream, but they additionally try to achieve two other goals: 1) output vertex or edge subsets as approximate solutions to the problem (not just real-valued estimates) and 2) use space that is sublinear in the number of edges or the number of vertices. Thus far, all previously known edge-differentially private algorithms for graph problems in the continual release setting do not meet the above benchmarks. Instead, they require computing exact graph statistics on the input [FichtenbergerHO21, Song18]. In this paper, we leverage sparsification to address the above shortcomings. Our edge-differentially private algorithms use sublinear space with respect to the number of edges in the graph while some also achieve sublinear space in the number of vertices in the graph. In addition, for most of our problems, we also output differentially private vertex subsets.
We make novel use of assorted sparsification techniques from the non-private streaming and static graph algorithms literature and achieve new results in the sublinear space, continual release setting for a variety of problems including densest subgraph, $k$-core decomposition, maximum matching, and vertex cover. In addition to our edge-differential privacy results, we use graph sparsification based on arboricity to obtain a set of results in the node-differential privacy setting, illustrating a new connection between sparsification and privacy beyond minimizing space. Previous node-differentially private algorithms only provide utility guarantees for the restrictive class of (nearly) bounded degree graphs [FichtenbergerHO21, Song18, jain2024time]. Our connection between sparsification via arboricity and node-privacy potentially opens up the space for additional results in both the static differential privacy and continual release models. We conclude with polynomial additive error lower bounds for edge-privacy in the fully dynamic setting.
@@ 1 Introduction
Today’s data is not only massive in size but also rapidly growing. As the world becomes increasingly digitized, the sensitive associations between individuals are becoming a key focus for data analysis. Such a focus also comes with privacy risks and the question of how to safeguard privacy. Differential privacy [DMNS06] is the gold standard for privacy protection, with a rich body of work focusing on static graphs [AU19, blocki2022privately, dinitz2024tight, DLRSSY22, ELRS22, kasiviswanathan2013analyzing, kalemaj2023node, LUZ24, mueller2022sok, NRS07, RS16, raskhodnikova2016differentially, Upadhyay13]. However, today’s graphs are not static; they are rapidly expanding, with new connections between individuals forming every minute. Streaming graph algorithms, a well-studied area in the non-private setting [assadi2022semi, AG11, EHW16, Esfandiari2018, feigenbaum2005graph, FMU22, HuangP19, henzinger1998computing, muthukrishnan2005data, mcgregor2014graph, mcgregor2015densest, mcgregor2018simple], are designed to handle such massively evolving datasets. In this model, updates to the dataset are received in a stream. Online streaming algorithms continuously release accurate graph statistics after each update [chen2023sublinear, halldorsson2016streaming, cormode2018approximating, GS24]; but these graph statistics may reveal private information.
@@
Our paper focuses on the well-known continual release model [DworkNPR10, CSS11] of differential privacy that promises a strong guarantee for online streaming outputs: the entire vector of all outputs satisfies differential privacy. The two notions of privacy we consider are (1) edge-differential privacy, where we require the output distributions of the algorithm on edge-neighboring stream inputs to be “close”, where edge-neighboring streams differ in one edge update, and, similarly, (2) node-differential privacy, where node-neighboring streams differ in all updates incident to one vertex. Despite the large body of differential privacy works on static graph algorithms, there are few works on graph continual release [FichtenbergerHO21, jain2024time, Song18] (and related graph models [upadhyay2021differentially]). While there has been substantial progress in graph continual release algorithms in recent years, there are some drawbacks to existing algorithms, including:
  • Being unable to release private (vertex subset) solutions in addition to their real-valued function outputs;
  • Requiring exact algorithms for computing graph statistics, thus necessitating linear space usage in the number of edges;
  • Only providing non-trivial node-privacy utility guarantees for the class of (nearly) bounded-degree graphs.
In this paper, we take the first steps to address these short-comings for insertion-only streams in the continual release model: we provide the first graph algorithms that return edge-differentially private vertex subset solutions (in addition to the value of the solutions) and use sublinear space for general graphs. The approximation guarantees of our algorithms nearly match the guarantees for their private static or non-private streaming counterparts. Moreover, as is standard in streaming, all of our results hold for a single pass over the stream. In the more challenging node-privacy setting, we derive algorithms with non-trivial utility guarantees for the class of bounded arboricity graphs, which are a strict super class of bounded degree graphs. [^1] Consider the star graph on $n$ vertices which has maximum degree $n-1$ but arboricity 1.
@@
Resolving these issues makes continual release graph algorithms useful alternatives to their non-private streaming counterparts for many applications. While non-graph private streaming literature has explored sublinear space [ChanLSX12, epasto2023differentially, upadhyay2019sublinear], this topic has not been considered for general graphs in continual release. In most use cases, streaming algorithms are used for datasets that are so large that algorithm designers assume that the entire dataset cannot fit in memory. Moreover, real-life applications often call for vertex subset solutions instead of just the values; for example, advertisers may want the members of the current densest community. Finally, it is preferable to protect the privacy of all of the connections of any user instead of any one connection for problems for which we can achieve non-trivial node-private solutions.
@@ 1.1 Sparsification
The main algorithmic technique we use to resolve all three of the aforementioned drawbacks is sparsification, traditionally used in the streaming, static, and dynamic graph algorithms literature. The idea behind sparsification is to create a smaller subgraph of the input graph, using a carefully chosen subset of edges from the stream, which (approximately) preserves the desired property that we want to return for our problem. Such sparsifiers exist for a large number of problems in the non-private literature, including for problems such as densest subgraph, $k$-core decomposition, maximum matching, and vertex cover. In general, graph sparsifiers are problem-specific, where each sparsifier provides guarantees for only one or two closely related problems. We design a systematic method of creating problem-specific differentially private sparsifiers for our problems, which we describe in more detail in the technical overview (Section 2).
Although several projection techniques for graph sparsification exist in the static, one-shot, and sliding window differential privacy literature [BBS13, CZ13, Day16, kalemaj2023node, kasiviswanathan2013analyzing, Upadhyay13, upadhyay2021differentially], such techniques are harder to use in the continual release setting, as the decision to remove a single edge can propagate temporally to many additional changes down the road.
Example 1.1.
Consider the naive sparsifier that removes all edges adjacent to nodes with degree greater than a public bound $\tilde{D}$. On edge-neighboring streams $S\sim S^{\prime}$ node $v$ may have degree $\tilde{D}$ in $S$ and degree $\tilde{D}+1$ in $S^{\prime}$. The naive sparsifier would remove all edges adjacent to $v$ in $S^{\prime}$ but not in $S$. Whereas the previous sensitivity of the two streams is $1$, the sparsified streams now have sensitivity $\tilde{D}$. Every release after the insertion of the last edge to $v$ in $S^{\prime}$ must use sensitivity $\tilde{D}$.
Hence, our techniques must preserve the edge edit distance (Definition 3.5) between neighboring streams. The first time-aware projection algorithms [^2] Defined to be projections that project a stream into one that is (nearly) degree-bounded without significantly increasing the edge edit distance. were designed very recently in [jain2024time] that lead to $\operatorname*{poly}(\log n)$ approximations for node-differential privacy when the input graph stream is guaranteed to have (nearly) bounded degree [FichtenbergerHO21, Song18]. However, it does not guarantee good approximations in more general graphs when $\omega(\operatorname*{poly}(\log(n)))$ vertices have degree $\omega(\operatorname*{poly}(\log(n)))$. We design a number of novel sparsification techniques that are distance-preserving [^3] This means that the sparsified outputs differ by no more than a constant number of edges on edge-neighboring streams and no more than $\alpha$ edges in node-neighboring streams where $\alpha$ is the arboricity. This concept is also referred to as stability in [jain2024time]. and can also be applied in the continual release setting. Our sparsification techniques directly lead to space savings for both edge and node-privacy for a variety of problems including densest subgraph, $k$-core decomposition, maximum matching, and vertex cover; as a bonus, we return vertex subsets for the densest subgraph problem, approximate core numbers for every vertex in core decomposition, and an implicit solution for vertex cover.
Using our sparsification techniques, we also demonstrate the first connection between sparsification and node-differentially private algorithms via arboricity. The class of bounded arboricity [^4] Arboricity is defined as the minimum number of forests to decompose the edges in a graph. A $n$ degree star has max degree $n-1$ and arboricity $1$. graphs is a more general class of graphs than bounded degree graphs; a simple example of a graph with large degree but small arboricity is a collection of stars. We broaden the scope of node differentially private algorithms for graph problems in the continual release model beyond bounded degree [FichtenbergerHO21, Song18] and close to bounded degree graph inputs [jain2024time]. We develop algorithms for maximum matching and vertex cover, where, given a public bound $\tilde{\alpha}$ on the arboricity of the graph, our algorithm returns error bounded by $\tilde{\alpha}$ when $\tilde{\alpha}$ upper bounds the maximum (private) arboricity of the stream. Our algorithms are always private, return good approximations when the public bound upper bounds the private graph property we require, and use sublinear space.
@@ 1.2 Our Contributions
Our paper gives the following set of results, stated informally here and given in more precise terms in their respective cited sections. We group our results by graph problems. Our upper bounds consider insertion-only streams where edge additions are given in the stream. Our algorithms produce $(\beta,\zeta)$-bicriteria approximations where $\beta$ is the multiplicative factor in the approximation and $\zeta$ is the additive error. Throughout the paper, we assume $n\leq T=O(\operatorname*{poly}(n))$. This is reasonable as $O(n^{2})$ (non-empty) updates are sufficient to have a complete graph. [^5] Our results can be modified to handle streams of length $T=\omega(\operatorname*{poly}(n)).$ We sometimes compare our results with static DP lower bounds because such lower bounds also hold in the insertion-only continual release setting. [^6] Otherwise, if we obtain better error in the continual release setting, we can solve the static problem with better error by inserting all of the edges of the static graph and taking the solution released at the end of the stream. A more in-depth discussion of related works can be found in Appendix A.
Maximum Matching (Section 4)
We give the following results for estimating the size of a maximum matching in bounded arboricity graphs. In this setting, $\tilde{\alpha}$ is a given public bound on the maximum arboricity $\alpha$ of the graph given by the input stream.777 Such public bounds are commonly assumed in non-private streaming literature [CJMM17, mcgregor2018simple]. We also briefly sketch how to remove this assumption in Section E.2 at the cost of more space and a worse approximation guarantee. See Table 1 for a comparison of our algorithms with previous results. We remark that the additive error of our algorithms nearly match the information-theoretic lower bound in bounded arboricity graphs up to logarithmic factors.
Firstly, in the truly sublinear model, where space usage is sublinear in the number of nodes in the graph, we give an edge-DP algorithm for estimating the size of a maximum matching in bounded arboricity graphs.
@@
Given a public bound $\tilde{\alpha}$, we obtain an $\varepsilon$-edge differentially private approximate maximum cardinality matching algorithm that outputs an approximate matching size in $O\left(\frac{\log^{2}(n)\log(\tilde{\alpha})}{\varepsilon}\right)$ space in the insertion-only continual release model. If $\tilde{\alpha}\geq\alpha$, where $\alpha$ is the (private) maximum arboricity of the input graph, then our algorithm returns a $\left((1+\eta)(2+\tilde{\alpha}),O\left(\frac{\log^{2}(n)}{\varepsilon}\right)\right)$-approximation of the size of the maximum matching.
We also derive the first maximum matching algorithms for the stronger model of node-DP beyond (nearly) bounded degree graphs.

### v3

# Sublinear Space Graph Algorithms in the Continual Release Model

@@ Abstract
The graph continual release model of differential privacy seeks to produce differentially private solutions to graph problems under a stream of edge updates where new private solutions are released after each update. Thus far, previously known edge-differentially private algorithms for most graph problems including densest subgraph and matchings in the continual release setting only output real-value estimates (not vertex subset solutions) and do not use sublinear space. Instead, they rely on computing exact graph statistics on the input [FichtenbergerHO21, Song18]. In this paper, we leverage sparsification to address the above shortcomings for edge-insertion streams. Our edge-differentially private algorithms use sublinear space with respect to the number of edges in the graph while some also achieve sublinear space in the number of vertices in the graph. In addition, for the densest subgraph problem, we also output edge-differentially private vertex subset solutions; no previous graph algorithms in the continual release model output such subsets.
We make novel use of assorted sparsification techniques from the non-private streaming and static graph algorithms literature to achieve new results in the sublinear space, continual release setting. This includes algorithms for densest subgraph, maximum matching, as well as the first continual release $k$-core decomposition algorithm. We also develop a novel sparse level data structure for $k$-core decomposition that may be of independent interest. To complement our insertion-only algorithms, we conclude with polynomial additive error lower bounds for edge-privacy in the fully dynamic setting, where only logarithmic lower bounds were previously known.
@@ 1 Introduction
Today’s data is not only massive in size but also rapidly growing. As the world becomes increasingly digitized, the sensitive associations between individuals are becoming a key focus for data analysis. Such a focus also comes with privacy risks and the question of how to safeguard privacy. Differential privacy [DMNS06] is the gold standard for privacy protection, with a rich body of work focusing on static graphs [AU19, blocki2022privately, dinitz2024tight, DLRSSY22, ELRS22, kasiviswanathan2013analyzing, kalemaj2023node, LUZ24, mueller2022sok, NRS07, RS16, raskhodnikova2016differentially, Upadhyay13]. However, today’s graphs are not static; they are rapidly expanding, with new connections between individuals forming every minute. Streaming graph algorithms, a well-studied area in the non-private setting [assadi2022semi, AG11, EHW16, Esfandiari2018, feigenbaum2005graph, FMU22, HuangP19, henzinger1998computing, muthukrishnan2005data, mcgregor2014graph, mcgregor2015densest, mcgregor2018simple], are designed to handle such massively evolving datasets. In this model, updates to the dataset are received in a stream. Online streaming algorithms continuously release accurate graph statistics after each update [chen2023sublinear, halldorsson2016streaming, cormode2018approximating, GS24], but these graph statistics may reveal private information.
@@
Our paper focuses on studying graph problems in the well-known continual release model [DworkNPR10, CSS11] of differential privacy that promises a strong guarantee for online streaming outputs: the entire vector of all outputs satisfies differential privacy. The notion of graph privacy we consider is edge-differential privacy, where we require the output distributions of the algorithm on edge-neighboring stream inputs to be “close”, where edge-neighboring streams differ in one edge update. Despite the large body of differential privacy works on static graph algorithms, there are few works on graph continual release [FichtenbergerHO21, jain2024time, Song18] and related graph models [upadhyay2021differentially]. While there has been substantial progress in graph continual release algorithms in recent years, there are some drawbacks to existing algorithms, including: 1) not releasing private (vertex subset) solutions in addition to their real-valued function outputs and 2) requiring exact algorithms for computing graph statistics, thus necessitating linear space usage in the number of edges.
In this paper, we take the first steps to address these shortcomings for insertion-only streams in the continual release model via the following contributions:
  • We give the first $k$-core decomposition algorithm in the continual release model.
  • We formulate the first continual release graph algorithms that return edge-differentially private vertex subset solutions for densest subgraph (in addition to the value of the solutions).
  • We give the first sublinear space algorithms for general graphs to solve densest subgraph, $k$-core decomposition, and matching problems.
The approximation guarantees of our algorithms nearly match the guarantees for their private static or non-private streaming counterparts. Moreover, as is standard in streaming, all of our results hold for a single pass over the stream.
@@
Resolving these issues closes the gap between continual release graph algorithms and their non-private streaming counterparts for many applications. While the private streaming literature has explored sublinear space outside the graph algorithms setting [ChanLSX12, epasto2023differentially, upadhyay2019sublinear, BlockiGMZ22], it remains unaddressed in general graphs for most problems in continual release. Furthermore, while streaming algorithms typically operate on datasets too large for memory, practical applications often require information about each vertex, not just aggregate values. For example, advertisers may want the members of the current densest community, not merely its density. Hence, achieving both goals–sublinear space combined with useful vertex information–is essential to the development of private graph algorithms.
@@ 1.1 Sparsification
The main algorithmic technique we use in this paper is sparsification, traditionally used in the streaming, static, and dynamic graph algorithms literature. Sparsification aims to simplify a large graph by strategically selecting a smaller set of its edges. This smaller “sketch” of the graph still captures the essential information needed for a specific problem. Such sparsifiers exist for a large number of problems in the non-private literature. In general, graph sparsifiers are problem-specific, where each sparsifier provides guarantees for only one or two closely related problems. We design novel methods for creating differentially private problem-specific graph sparsifiers in continual release, which we describe in more detail in the technical overview (Section 2).
Using sparsification techniques, we design a number of new continual release algorithms that achieve the same utility guarantees as previous algorithms for the problem in the static DP and non-private streaming settings up to logarithmic factors. Our sparsification techniques directly lead to space savings in the edge-privacy setting for a variety of problems including densest subgraph, $k$-core decomposition, and maximum matching. We return vertex subsets for the densest subgraph problem, approximate $k$-core numbers for every vertex in core decomposition, and estimates for the size of the maximum matching, all using sublinear space in the number of edges. For maximum matching, we even use space that is sublinear in the number of vertices. [^1] While all non-private streaming algorithms use space that is sublinear in the number of total edges, very few use space sublinear in the number of vertices, although results sublinear in the number of vertices are most desirable.
@@ 1.2 Our Contributions
Our paper gives the following set of results, stated informally here and given in more precise terms in their respective sections. We group our results by graph problems for $n$-vertex $m$-edge graph streams over $T$ time steps and our multiplicative guarantees are given in terms of a constant factor $\eta>0$.
For our algorithmic results, we consider insertion-only streams where (possibly empty) edge insertions are given in the stream. We may assume without loss of generality that each non-empty edge is inserted at most once, as we can treat an insertion operation for an existing edge as an empty insertion. Without requiring privacy, this closely resembles the well-studied online streaming setting (see related works in Section 3). Our algorithms also incur logarithmic additive noise, which is necessary, as demonstrated by lower bounds that hold even in the static model (see Section 1.2).
Our lower bounds are given for fully dynamic streams where edges can be inserted and deleted multiple times, also known as the turnstile model. In this setting, we show that polynomial additive error is necessary, demonstrating a fundamental difficulty in maintaining small errors in the presence of both insertions and deletions.
@@
All of our results in the continual release model are given for edge-neighboring streams where the two streams differ by one edge update; results in the static DP setting are given for edge-neighbors (graphs that differ by one edge). Throughout the paper, we will use the phrase $\varepsilon$-edge DP to refer to these settings. Our algorithms produce $(\beta,\zeta)$-bicriteria approximations where $\beta$ is the multiplicative factor in the approximation and $\zeta$ is the additive error. Our approximation bounds hold, with high probability, for the released solution after every update in the stream. Throughout the paper, we assume $n\leq T=O(\operatorname*{poly}(n))$. This is reasonable as $O(n^{2})$ (non-empty) updates are sufficient to have a complete graph. [^2] Our results can be modified to handle streams of length $T=\omega(\operatorname*{poly}(n)).$ We compare our results with static edge-DP lower bounds as such lower bounds also hold in the insertion-only continual release setting. [^3] Otherwise, if we obtain better error in the continual release setting, we can solve the static problem with better error by inserting all of the edges of the static graph and taking the solution released at the end of the stream. A more in-depth discussion of related works can be found in Section 3. Our entire set of results can be found in Section 1.2, including baseline algorithms and lower bounds in the static edge-DP and non-private streaming settings.
Table 1: Our results are highlighted in green. Non-private bounds are shown in gray. All results are for edge-privacy. Bicriteria approximations $(\beta,\zeta)$ are given, where $\beta$ is the multiplicative factor and $\zeta$ is the additive error. $\tilde{\alpha}$ is a public upper bound on the maximum arboricity of the stream; our approximation guarantees for maximum matching hold when $\tilde{\alpha}$ upper bounds the private arboricity and our privacy guarantees always hold. $\alpha$ in the non-private streaming bound is the maximum arboricity in the stream.
{NiceTabular}
—c—c—c—c—c—c— Problem Space Approximation Reference Model Solution Type
$k$-Core Decomp. $\tilde{O}\left(\frac{n}{{\eta^{4}}\varepsilon}\right)$ $\left(2+\eta,O\left(\frac{\log^{3}(n)}{{\eta^{2}}\varepsilon}\right)\right)$ Theorem 5.1 Insertion Only Upper Bound
$\Theta(m+n)$ $\left(1,O\left(\frac{\log(n)}{\varepsilon}\right)\right)$ [DLL23] Static Upper Bound
$\left(\beta,\Omega\left(\frac{\log n}{\varepsilon\beta}\right)\right)$ [HSZ24] Static Lower bound
$\tilde{O}\left(\frac{n}{\eta^{2}}\right)$ $(1+\eta,0)$ [Esfandiari2018] Non- Private Streaming Upper Bound
Densest Subgraph $\tilde{O}\left(\frac{n}{{\eta^{2}}\varepsilon}\right)$ $\left(2+\eta,O\left(\frac{\log^{2}(n)}{{\eta}\varepsilon}\right)\right)$ Theorem 6.1 Insertion Only Upper Bound Vertex Subset
$\tilde{O}\left(\frac{n}{{\eta^{5}}\varepsilon}\right)$ $\left(1+\eta,O\left(\frac{\log^{5}(n)}{{\eta^{4}}\varepsilon}\right)\right)$ Theorem 6.2 Insertion Only Upper Bound Vertex Subset
$\Theta(m+n)$ $\left(1+\eta,O\left(\frac{\log^{4}(n)}{\eta^{3}\varepsilon}\right)\right)$ [DLRSSY22] Static Upper Bound Vertex Subset
$\Theta(m+n)$ $\left(2,O\left(\frac{\log(n)}{\varepsilon}\right)\right)$ [DLL23] Static Upper Bound Vertex Subset
$\Theta(m+n)$ $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$ [dinitz2024tight] Static Upper Bound Vertex Subset
$\left(\beta,\Omega\left(\frac{1}{\beta}\sqrt{\frac{\log(n)}{\varepsilon}}\right)\right)$ [AHS21] Static Lower bound
$\Theta(m)$ $\left(1+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$ [FichtenbergerHO21] Insertion Only Upper Bound Density Only
$\tilde{O}\left(\frac{n}{\eta^{2}}\right)$ $(1+\eta,0)$ [mcgregor2015densest] [EHW16] Non- Private Streaming Upper Bound
Maximum Matching Size $O\left(\frac{\log^{2}(n)\log(\tilde{\alpha})}{{\eta^{2}}\varepsilon}\right)$ $\left((1+\eta)(2+\tilde{\alpha}),O\left(\frac{\log^{2}(n)}{{\eta}\varepsilon}\right)\right)$ Theorem 7.1 Insertion Only Upper Bound Arboricity
$\Theta(m)$ $\left(1+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$ [FichtenbergerHO21] Insertion Only Upper Bound
$\left(1,\Omega\left(\log(T)\right)\right)$ [FichtenbergerHO21] Insertion Only Lower bound
$\left(1,\Omega\left(\min\left(\sqrt{\frac{n}{\varepsilon}},\frac{T^{1/4}}{\varepsilon^{3/4}},n,T\right)\right)\right)$ Theorem 8.1 Fully Dynamic Lower Bound
$\left(1,\Omega\left(\log(T)\right)\right)$ [FichtenbergerHO21] Fully Dynamic Lower Bound
$O\left(\frac{\log(n)\log(\alpha)}{\eta^{2}}\right)$ $\left((1+\eta)(2+\alpha),0\right)$ [mcgregor2018simple] Non- Private Streaming Upper Bound Arboricity
Triangle Count $\left(1,\Omega\left(\min\left(\sqrt{\frac{n}{\varepsilon}},\frac{T^{1/4}}{\varepsilon^{3/4}},n,T\right)\right)\right)$ Section 8.1 Fully Dynamic Lower Bound
$\left(1,\Omega\left(\log(T)\right)\right)$ [FichtenbergerHO21] Fully Dynamic Lower Bound
Connected Components Counting $\left(1,\Omega\left(\min\left(\sqrt{\frac{n}{\varepsilon}},\frac{T^{1/4}}{\varepsilon^{3/4}},n,T\right)\right)\right)$ Theorem 8.2 Fully Dynamic Lower Bound
$k$-Core Decomposition (Section 5). We show the first $\varepsilon$-edge DP algorithm for the $k$-core decomposition problem in the continual release model. Our algorithm returns approximate $k$-core numbers at every timestep while using $o(m)$ space. The additive errors of our algorithms are near-optimal as they match the recent additive error lower bounds up to logarithmic factors (see Section 1.2). We note that these lower bounds hold even when allowing for constant multiplicative error.
@@
We obtain a $\left(2+\eta,O\left(\frac{\log^{3}(n)}{{\eta^{2}}\varepsilon}\right)\right)$-approximate $\varepsilon$-edge differentially private $k$-core decomposition algorithm that outputs core number estimates in the insertion-only continual release model using $\tilde{O}\left(\frac{n}{\eta^{4}\varepsilon}\right)$ space.
The essence of our $k$-core algorithm is a novel sparse level data structure that may be of independent interest beyond our work. To the best of our knowledge, such a data structure did not previously exist in the privacy or graph algorithms literature.
Densest Subgraph (Section 6). We obtain the first $\varepsilon$-edge DP densest subgraph algorithms that return subsets of vertices with near-optimal density in the continual release model. Our algorithms also use $o(m)$ space. We emphasize that the approximation guarantees nearly match, up to logarithmic terms, the guarantees of their private static or non-private streaming counterparts. In turn, these nearly match the information-theoretic lower bound up to logarithmic factors. Again, we emphasize that these lower bounds hold even when allowing for constant multiplicative error.